Use the worked example above to help you solve thi
Answer #1

(1)

del PE = -q Ex del x

= - (1.6 * 10^-19) ( 1.50 * 10^3 )( 0.0496 m - ( -0.0290)

=-1.88 * 10^-17 J

(2)

del PE = -q Ex del x

= - (-1.6 * 10^-19) ( 1.50 * 10^3 )( 0.13 m - ( -0.0290)

=3.81 * 10^-17 J

(3)

del PE = -q Ex del x

= - (-1.6 * 10^-19) (- 1.50 * 10^3 )( 0.072m - ( 0.0380)

=-8.16 * 10^-18 J